# Contextuality of the probability current in quantum mechanics

### v1

@@ Abstract
We revisit an argument proposed by Hardy [1] concerning local realistic theories, but in terms of probability currents in standard quantum mechanics and of the trajectories obtained from its flux lines. We emphasize surprising properties of the trajectories in configuration space, in particular their (quasi) discontinuous variations when the context (the experimental setup) is changed. This occurs in both Galilean and Einsteinian relativity. In the latter case, discontinuous variations also appear without changing the setup, when the Lorentzian reference frame is changed.
The purpose of this article is to discuss the properties of the probability current in relativistic quantum mechanics. Within Galilean relativity, one can use the Schrödinger equation, which implies the existence of a local conservation law of the probability containing a current $\bm{J}(\bm{r}_{1},\bm{r}_{2},\ldots,\bm{r}_{N},)$; this current is defined in the configuration space of the physical system (the space associated with the $N$ spatial coordinates of its particles). One can then build trajectories, which are at each point tangent to $\bm{J}$. When a different Galilean reference frame is used, the components of $\bm{J}$ are transformed as those of the current of a classical fluid. It is then possible to develop pictorial views of the motion of the probability fluid and, if one wishes to introduce additional variables in quantum mechanics, to build the dBB (de Broglie-Bohm) theory [2, 3, 4]. Within Einsteinian relativity, we will see that the situation is more complex, and that it is not possible to define a probability current that takes consistent values in different Lorentz frames.
The essence of our discussion is very similar to that already given by Hardy [1], but here we emphasize the role of the probability current, instead of the “elements of reality”. In § 1, we first analyze the experiment within Galilean relativity, with absolute time. We show that, even in this simple case, the probability current exhibits unexpected contextual properties; in particular, it can change abruptly when the experimental conditions (the context) are slightly modified. In § 2 we discuss the same experiment within Einsteinian relativity, and show that the properties of the probability current are even more surprising: when the same experiment is described in a different Lorentz reference frame, the probability current may look very different; it may even exhibit (almost) discontinuous variations that have very little to do with those of the current of a classical fluid. This puts severe constraints on the construction of theories with additional variables.
@@ 1 Galilean relativity
Following Hardy [1], we consider two interferometers $\mathcal{A}$ and $\mathcal{B}$ (figure 1). The particles crossing these devices are entangled since, if the particles fly across the arms $a_{2}$ and $b_{2}$ in the middle of the figure, they interact and annihilate each other by some process.
Figure 1: Each of the two interferometers $\mathcal{A}$ and $\mathcal{B}$ receives a particle. The oval symbolizes an interaction process that destroys the two particles if they both take the middle paths. This process creates an entanglement between the two particles.
We denote by $|a_{0}\rangle$ and $|b_{0}\rangle$ the states of the particles entering the interferometers, by $|a_{1}\rangle$ and $|a_{2}\rangle$ (first particle), $|b_{1}\rangle$ and $|b_{2}\rangle$ (second particle), their quantum states when they propagate inside the inteferometers, as indicated in figure 1. Finally $|A_{1}\rangle$ and $|A_{2}\rangle$, $|B_{1}\rangle$ and $|B_{2}\rangle$ are the quantum states of each particle after it crosses the output beam splitter. The effect of these output beam splitters is described by the substitutions [7]:
@@ 1.1 State vectors
Just after crossing the two first beam splitters, the system is in state:
@@
1.2 Probability currents and trajectories
@@
We now consider situations where two particles, described by wave packets, enter simultaneously the two interferometers $\mathcal{A}$ and $\mathcal{B}$. The Schrödinger probability current $\bm{J}$ has 6 components in configuration space (3 containing the derivatives of the wave function with respect to the coordinates of the first particle, 3 those for the other particle), and provides the direction along which the pair of the two wave packets moves in this space. One can then build continuous trajectories in configuration space that are tangent to $\bm{J}$ at each of their points. Since they are nothing but pairs of trajectories in usual 3D space, and in order to avoid confusion with ordinary trajectories, we will call them “bi-trajectories”. In general (if the wave function is a not product of single particle wave functions), the current depends on the positions of both particles; the 3D trajectory of each particle depends on the position of the other on its own trajectory.
@@
The properties of bi-trajectories may look significantly different from those of a 3D trajectory for a single particle. While in the latter case, a 3D no-crossing rule limits the possible patterns, for bi-trajectories the no-crossing rule applies only in the configuration space, so that crossing in ordinary space are indeed possible. For instance, several different trajectories my start from the same point in interferometer $\mathcal{A}$, provided they are associated with different trajectories in the other interferometer. Moreover, the direction of each 3D trajectory may be affected by the insertion of a beam splitter recombining two wave functions in the other interferometer. Of course these properties occur only in the presence of entanglement, which introduces quantum nonlocal effects.
@@
With wave packets, the probabilities of two-particle measurement results can be obtained as the time integral of the 6D flux of the probability current through the input surface of a detector. This operation should nevertheless be performed only with the values of the current at a given time; one should not use the values of the unperturbed probability current at different times to obtain correlation functions for measurement at different times. To correctly obtain these functions, one must take into account the effect of the first measurement on the state vector, which changes the probability current in both interferometers (even if the measurement is performed in only one of them). In Bohmian theory for instance, the calculation of time correlation functions requires that the perturbation of the first measurement on the evolution of the Bohmian positions is included, before the probability of the second measurement is calculated [9].
Inside the interferometers, the possible trajectories associated with the setup shown in figure 1 are obtained from the state vector of equation (3), while the state written in (5) determines the trajectories after the particles have crossed the output beam splitters. No pair of trajectories follows the path $a_{2}b_{2}$, since the corresponding component of the state vector vanishes: if one 3D trajectory takes the inner path, the other has to take the outer path. Whatever the detailed configurations of the bi-trajectories is, relation (5) indicates that they reach the 4 output channels, since they indicate the existence of a probability current that brings the probability into these regions.
@@ 1.3 Different configurations
Now, consider the situation shown in figure 2, where the output of beam splitter of $\mathcal{B}$ has been pushed to the right. The particle in interferometer $\mathcal{A}$ then crosses its output beam splitter before the other; in the intermediate time, the state of the system is given by relation (4a). This expression shows that, if the trajectory is such that the particle in interferometer $\mathcal{B}$ follows path $b_{1}$, the particle in interferometer $\mathcal{A}$ necessarily exits through $A_{1}$. The only the bi-trajectories where one particle reaches $A_{2}$ must have the other following path $b_{2}$.
@@ 1.4 Discussion
According to relation (3) the two particles cannot be found simultaneously in $a_{2}$ and $b_{2}$, which shows that no bi-trajectory takes this path. Relation (5) indicates that some bi-trajectories reach $A_{2}$ and $B_{2}$ (more precisely, 1/12 of the trajectories). But, in § 1.3, we concluded in one case that the only trajectories reaching $A_{2}$ have to pass trough $b_{2}$, and in the other case that the only trajectories reaching $B_{2}$ have to pass through $a_{2}$; combined together, these two statements seem to indicate that the trajectories reaching $A_{2}$ and $B_{2}$ have to pass through $a_{2}$ and $b_{2}$, which is impossible as we just saw.
@@
The mechanism behind this sensitivity of the probability current to the whole experimental setup is nonlocal in ordinary 3D space. When the particle in $\mathcal{B}$ crosses its beam splitter, components of the wave function that were spatially disjoint (in terms of the positions of that particle) are made to overlap, which introduces interference terms for the wave function guiding the particle in $\mathcal{A}$. This nonlocal effect is directly seen on the expressions (4a) to (5) of the quantum state of the physical system.
An striking illustration of this nonlocality of the probability current is the appearance of discontinuities when the experimental situations are modified, more precisely very rapid changes of the current lines when one element of the experimental device is slightly changed. When the position of one beam splitter is moved in order to progressively change the setup from that of figure 2 to that of figure 3, the probability current undergoes a sudden variation when the situation of figure 2 is crossed. Initially, all bi-trajectories reaching $A_{2}$ and $B_{2}$ pass through $a_{1}$ and $b_{2}$, while at the crossing they jump discontinuously to $a_{2}$ and $b_{1}$ (in dBB theory, the evolution of the Bohmian position necessarily exhibits the same discontinuity).
More precisely, when wave packets of finite length $\ell$ are crossing the output beam splitters, the state vector may actually be a superposition of states of the form (4) and (5). The rapid change of the pattern of the current lines should then take place when the length $L$ of one interferometer varies by $\ell$ around the length of the other interferometer. So, in fact no real mathematical discontinuity should occur, but just a very rapid change between two regimes of the current lines for a small variation of $\ell$. Since $L$ may be arbitrarily large, the crossover may seen as a quasi-discontinuity at the scale of $L$.
Hardy has also discussed this situation in terms of “empty waves” and their nonlocal effects. In any theory where localized particles are guided by a wave, the very possibility of observing pairs of particles exiting through outputs $A_{2}$ and $B_{2}$ is indeed paradoxical: on one hand, such an event is possible only if the particles interact by taking the paths $a_{2}$ and $b_{2}$, as seen by comparing relations (5) and (6); but, on the other hand, they are supposed to disappear if they follow these trajectories !
Ref. [10] discusses another version of the experiment, where the oval in figure 1 symbolizes a dephasing interaction, instead of the mutual annihilation of the two particles. This variant of the experiment may be relevant in the context of the detection of quantum effects induced by gravity [11]. Interestingly, the two versions of the experiment lead to very similar conclusions.
@@ 2.1 General properties
Within a relativistic quantum theory, we consider a physical system made of a fixed number $N$ of massive particles, and we assume that the theory contains a local conservation law of the probability. This law involves a probability current defined in configuration space, taking different value in different Lorentz frames; nevertheless, in each frame, a local law of conservation of the probability can be written with its own current. This current is closely related to the probabilities of measurement results since, with wave packets, the flux of this current (integrated over time) through appropriate surfaces (for instance the input surfaces of detectors) allows one to calculate probabilities, i.e. relativistic invariants.
For a system of two particles, the flux has to be calculated at a given time through a surface in a 6D space, in other words through a pair or surfaces in ordinary 3D space. As above, the probability current $\bm{J}$ allows one to define bi-trajectories, which are tangent to $\bm{J}$ at each point of configuration space. One can then obtain the probabilities of measurement results by making statistics over the trajectories, which are nothing but statistics over their starting point (with a weight given by the initial quantum probability of presence). This provides a view of the theory where the random character does not appear during measurements, but only in the value of the initial positions of the trajectories.
The current can be used in each Lorentzian frame to obtain the probability of any measurement performed at a given time in this frame. We will therefore use it for simultaneous measurements performed with two detectors sitting at any place in each interferometer (before or after the beam splitters), but not to derive correlations between results of measurements performed at different times – since this would require to take into account the perturbation of the first measurement on the current [9]. This necessity can be seen as a consequence of the von Neumann projection postulate, or of the entanglement of the system with the measurement apparatus (for instance, within dBB theory, the Bohmian positions within this apparatus play an important role [12]).
@@ 2.2 Different views in different Lorentzian frames
We now consider an double interference experiment which, in the Lorentz frame of the laboratory $\mathcal{L}_{0}$, corresponds to the scheme shown in figure 1. In another Lorentz frame, the same experiment is seen differently. For instance, figure 4 shows a time-space diagram describing the experiment in a frame $\mathcal{L}_{1}$ moving with a positive velocity $v$ along axis $Ox$. The output beam splitter is then crossed in interferometer $\mathcal{A}$ before than it is in interferometer $\mathcal{B}$. Thus, in this Lorentz frame, one can consider a simultaneous measurement of the output channel in interferometer $\mathcal{A}$ while the other particle is still in interferometer $\mathcal{B}$; the probability of any result can then be obtained by time integration of the current in this frame. To obtain this probability, we consider the same measurement performed in the laboratory frame, which becomes an time-shifted measurement experiment and corresponds to the configuration of figure 2. The probability of this two time measurement in the laboratory frame is then obtained by using the standard Schrödinger theory and the expression (4a) of the state vector. Finally, we use the fact that probabilities are relativistic scalars to obtain the probability of the measurement as seen in the moving frame $\mathcal{L}_{1}$, and from this infer properties of the probability current and trajectories in this frame.
For the same reason as in § 1, if the detectors are inserted in channel $A_{2}$ and $b_{1}$, the probability vanishes, but does not if they are inserted in channel $A_{2}$ and $b_{2}$. All bi-trajectories that reach $A_{2}$ therefore have to pass through $b_{2}$. If we extend the (unperturbed) trajectories by continuity we see that, in Lorentz frame $\mathcal{L}_{1}$, all bi-trajectories reaching outputs $A_{2}$ and $B_{2}$ necessarily pass through $b_{2}$. Since, as above, expression (3) shows that no trajectory passes through channels $a_{2}$ and $b_{2}$, we conclude that these trajectories necessarily pass through $a_{1}$ and $b_{2}$.
@@
Figure 4: This figure shows the experiment of figure 1 in another Lorentz frame, moving with a positive velocity $v$ in the direction of the vertical $Ox$ axis in the figure. The particles are assumed to have a non-zero mass. In interferometer $\mathcal{A}$, in some sections the propagation is parallel to $Ox$ to axis, which increases the velocity; in interferometer $\mathcal{B}$ it is antiparallel, which decreases the velocity. As a consequence, the crossing of the output beam splitter occurs first in interferometer $\mathcal{A}$ as in figure 2.

### v2

@@ Abstract
We revisit an argument proposed by Hardy [1] concerning local realistic theories, but in terms of the motion of the probability fluid and its probability current within standard quantum mechanic. We emphasize surprising properties of the flux lines of this current in configuration space, in particular (quasi) discontinuous variations when the context (the experimental setup) is changed. This occurs in both Galilean and Einsteinian relativity. In the latter case, discontinuous variations also appear without changing the setup, when the Lorentzian reference frame is changed. A Lorentz invariant definition of the motion of a probability fluid is therefore impossible.
The purpose of this article is to discuss the properties of the probability current in relativistic quantum mechanics. Within Galilean relativity, one can use the Schrödinger equation, which implies the existence of a local conservation law of the probability containing a current $\bm{J}(\bm{r}_{1},\bm{r}_{2},\ldots,\bm{r}_{N},)$; this current is defined in the configuration space of the physical system (the space associated with the $N$ spatial coordinates of its particles). One can then build lines of flux that are at each point tangent to $\bm{J}$. When a different Galilean reference frame is used, the components of $\bm{J}$ are transformed as those of the current of a classical fluid. It is then possible to develop pictorial views of the motion of the probability fluid and, if one wishes to introduce additional variables in quantum mechanics, to build the dBB (de Broglie-Bohm) theory [2, 3, 4] by adding positions variables. Nevertheless, in this article, we remain within standard quantum mechanics and limit ourselves to the properties of the usual probability current. Within Einsteinian relativity, we will see that the situation is more complex, and that it is not possible to define a probability current that takes consistent values in different Lorentz frames.
The essence of our discussion is very similar to that already given by Hardy [1], but here we emphasize the motion of the probability fuild and the role of the probability current, instead of the “elements of reality”. In § 1, we first analyze the experiment within Galilean relativity, with absolute time. We show that, even in this simple case, the probability current exhibits unexpected contextual properties; in particular, it can change abruptly when the experimental conditions (the context) are slightly modified. In § 2 we discuss the same experiment within Einsteinian relativity, and show that the properties of the probability current are even more surprising: when the same experiment is described in a different Lorentz reference frame, the probability current may look very different; it may even exhibit (almost) discontinuous variations that have very little to do with those of the current of a classical fluid. This puts severe constraints on the construction of theories with additional variables [5].
@@ 1 Galilean relativity
Following Hardy [1], we consider two interferometers $\mathcal{A}$ and $\mathcal{B}$ (figure 1). If the particles fly across the arms $a_{2}$ and $b_{2}$ in the middle of the figure, they interact and annihilate each other by some process. The two-particle system, if it survives, can only propagate in the three other paths $a_{1}b_{1}$, $a_{2}b_{1}$ , or $a_{1}b_{2}$, and becomes entangled.
Figure 1: Each of the two interferometers $\mathcal{A}$ and $\mathcal{B}$ receives a particle. The oval symbolizes an interaction process that annihilates the two particles if they both take the middle paths. To survive, the system of two particles has to propagate in the other combinations of paths, which results in quantum entanglement.
We denote by $|a_{0}\rangle$ and $|b_{0}\rangle$ the quantum states of the particles entering the interferometers. These states can be seen, either as plane waves of well defined energy-momentum, or, in a second step, as wave packets propagating in time and obtained by linearly combining these plane waves. Similarly, we denote by $|a_{1}\rangle$ and $|a_{2}\rangle$ (first particle), $|b_{1}\rangle$ and $|b_{2}\rangle$ (second particle), their quantum states when they propagate inside the inteferometers, as indicated in figure 1. Finally $|A_{1}\rangle$ and $|A_{2}\rangle$, $|B_{1}\rangle$ and $|B_{2}\rangle$ are the quantum states of each particle after it crosses the output beam splitter.
The effect of these output beam splitters is described by the substitutions [8]:
@@ 1.1 State vectors
Just after crossing the two first beam splitters, the system is in state [^1] As usual in this kind of discussion, for instance in Ref. [1] or in volume III of the Feynman lectures of physics, the kets written in the equations can be seen in two ways: either plane waves of well defined momentum, or wave packets obtained by linearly combining these plane waves. Explicitly writing wave packets would make the equations more complicated, but not change anything in the reasonings.:
@@
(since we are interested only in systems that have survived the annihilation process, we have normalized this new state to unity).
@@
1.2 Probability currents and bi-trajectories
We now consider situations where two particles, described by wave packets, enter simultaneously the two interferometers $\mathcal{A}$ and $\mathcal{B}$. The probability current in configuration space $\bm{J}(\bm{r}_{1},\bm{r}_{2})$ associated with a Schrödinger wave function $\psi(\bm{r}_{1},\bm{r}_{2})$ is [^2] We assume here that the particles have a non-zero mass, as is the case for atomic interferometry experiments (in this case, beam splitters can be obtained with non-resonant lasers and interference gratings). If the particles are polarized photons, a different expression of the current should be used.:
[…]
@@ 1.2 Probability currents and bi-trajectories
where c.c. indicates a complex conjugate. The nabla derivation operator has 6 components in configuration space (3 containing the derivatives of the wave function with respect to the coordinates of the first particle, 3 those for the other particle), and provides the direction along which the pair of the two wave packets moves in this space. One can then build continuous current lines (or flux lines) in configuration space that are tangent to $\bm{J}$ at each of their points. Since they are nothing but pairs of lines in usual 3D space, and since in dBB theory each of these lines correspond to a trajectory, we will call them “bi-trajectories”. Nevertheless our discussion remains at a purely wave description, without introducing Bohmian positions of the two particles (the variables $\bm{r}_{1}$ and $\bm{r}_{2}$ are just coordinates labelling the points in configuration space), so that this is just a convenient convention of language.
@@
In general (if the wave function is a not product of single particle wave functions), the current is not a product of functions of $\bm{r}_{1}$ and $\bm{r}_{2}$; the 3D current line in each interferometer depends on the current line in the other. We note that the properties of bi-trajectories may look significantly different from those of a 3D current line for a single particle. While in the latter case, a 3D no-crossing rule limits the possible patterns [4], for bi-trajectories the no-crossing rule applies only in the configuration space, so that crossings in ordinary space are indeed possible. For instance, several different current lines may start from the same point in interferometer $\mathcal{A}$, provided they are associated with different lines in the other interferometer. Similarly, the direction of each 3D current line may be affected by the insertion of a beam splitter recombining two wave functions in the other interferometer. Of course these mutual dependences occur only in the presence of entanglement, which introduces quantum nonlocal effects.
@@
With wave packets, the probabilities of two-particle measurement results can be obtained as the time integral of the 6D flux of the probability current through the input surface of a detector [^3] This operation should nevertheless be performed only with the values of the current at a given time; one should not try to use the values of the unperturbed probability current at different times to obtain correlation functions for measurement at different times. A correct calculation requires that one should take into account the effect of the first measurement on the state vector, which changes the probability current in both interferometers (even if the measurement is performed in only one of them). In Bohmian theory for instance, the calculation of time correlation functions requires that the perturbation of the first measurement on the evolution of the Bohmian positions is included, before the probability of the second measurement is calculated [9].. Inside the interferometers, the possible bi-trajectories associated with the setup shown in figure 1 are obtained from the state vector of equation (3), while the state written in (5) determines them after the particles have crossed the output beam splitters. No bi-trajectory follows the path $a_{2}b_{2}$, since the corresponding component of the state vector vanishes: if one 3D current line takes the inner path, the other has to take the outer path. Whatever the detailed configurations of the bi-trajectories is, relation (5) indicates that they always reach the 4 output channels, since they indicate the existence of a probability current that brings the probability into these regions.
@@ 1.3 Different configurations
Now, consider the situation shown in figure 2, where the output of beam splitter of $\mathcal{B}$ has been pushed to the right. The particle in interferometer $\mathcal{A}$ then crosses its output beam splitter before the other; in the intermediate time, the state of the system is given by relation (4a). This expression shows that, if the bi-trajectory is such that the particle in interferometer $\mathcal{B}$ follows path $b_{1}$, the particle in interferometer $\mathcal{A}$ necessarily exits through $A_{1}$. The only bi-trajectories where one particle reaches $A_{2}$ must have the other following path $b_{2}$.
@@ 1.4 Discussion
According to relation (3) the two particles cannot be found simultaneously in $a_{2}$ and $b_{2}$, which shows that no bi-trajectory takes this path. Relation (5) indicates that some bi-trajectories reach $A_{2}$ and $B_{2}$ (more precisely, 1/12 of them). But, in § 1.3, we concluded in one case that the only bi-trajectories reaching $A_{2}$ have to pass trough $b_{2}$, and in the other case that the only bi-trajectories reaching $B_{2}$ have to pass through $a_{2}$; combined together, these two statements seem to indicate that the bi-trajectories reaching $A_{2}$ and $B_{2}$ have to pass through $a_{2}$ and $b_{2}$, which is impossible as we just saw.
@@
The mechanism behind this sensitivity of the probability current to the whole experimental setup is nonlocal in ordinary 3D space. When the particle in $\mathcal{B}$ crosses its beam splitter, components of the wave function that were spatially disjoint (in terms of the positions of that particle) are made to overlap, which introduces interference terms for the wave function guiding the probability current line in $\mathcal{A}$. This nonlocal effect is directly seen on the expressions (4a) to (5) of the quantum state of the physical system.
A striking illustration of this nonlocality of the probability current is the appearance of discontinuities when the experimental situations are modified, more precisely very rapid changes of the current lines when one element of the experimental device is slightly changed. When the position of one beam splitter is moved in order to progressively change the setup from that of figure 2 to that of figure 3, the probability current lines undergo a sudden variation when the situation of figure 2 is crossed. Initially, all bi-trajectories reaching $A_{2}$ and $B_{2}$ pass through $a_{1}$ and $b_{2}$, while at the crossing they jump discontinuously to $a_{2}$ and $b_{1}$ (in dBB theory, the evolution of the Bohmian position necessarily exhibits the same discontinuity).
More precisely, while wave packets of finite length $\ell$ are crossing the output beam splitters, the state vector may actually be a superposition of states of the form (4) and (5). The rapid change of the pattern of the current lines should then take place when the length $L$ of one interferometer varies by $\ell$; in fact no real mathematical discontinuity should then occur, but just a very rapid change between two regimes of the current lines for a small variation of $\ell$. At the scale of $L$, which may be arbitrarily large with respect to $\ell$, the crossover may appear as a quasi-discontinuity.
Hardy has also discussed this situation in terms of “empty waves” and their nonlocal effects [10]. In any theory where localized particles are guided by a wave, the very possibility of observing pairs of particles exiting through outputs $A_{2}$ and $B_{2}$ is indeed surprising: on one hand, such an event is possible only if the particles interact by taking the paths $a_{2}$ and $b_{2}$, as seen by comparing relations (5) and (6); but, on the other hand, they are supposed to disappear if they follow this bi-trajectory !
Ref. [11] discusses another version of the experiment, where the oval in figure 1 symbolizes a dephasing interaction, instead of the mutual annihilation of the two particles. This variant of the experiment may be relevant in the context of the detection of quantum effects induced by gravity [12]. Simple calcualtions show that the two versions of the experiment, annihilation or dephasing, lead to very similar conclusions.
@@ 2.1 General properties
Within a relativistic quantum theory, we consider a physical system made of a fixed number $N$ of massive particles, and we assume that the theory contains a local conservation law of the probability, involving a $N$-particle probability current defined in configuration space. For example Refs. [13] and [5]) propose expressions of the probability current for relativistic Dirac particles that may be entangled but do not interact, which is sufficient for our discussion. Of course, the current takes different values in different Lorentz frames but, in each frame, a local law of conservation of the probability can be written with its own current. This current is necessarily related to the probabilities of measurement results since, with wave packets, the flux of this current (integrated over time) through appropriate surfaces (for instance the input surfaces of detectors) allows one to calculate probabilities, i.e. relativistic invariants.
For a system of two particles, the flux has to be calculated at a given time through a surface in a 6D space, in other words through a pair or surfaces in ordinary 3D space. As above, the probability current $\bm{J}$ allows one to define bi-trajectories, which are tangent to $\bm{J}$ at each point of configuration space. One can then obtain the probabilities of measurement results by making statistics over the bi-trajectories, which are nothing but statistics over their starting point (with a weight given by the initial quantum probability of presence). This provides a view of the theory where the random character does not appear during measurements, but only in the value of the initial positions of the bi-trajectories.
The current $\bm{J}$ can be used in each Lorentzian frame to obtain the probability of any measurement performed at a given time in this frame. We will therefore use it for simultaneous measurements performed with two detectors sitting at any place in each interferometer (before or after the beam splitters), but not to derive correlations between results of measurements performed at different times – since this would require to take into account the perturbation of the first measurement on the current [9]. This necessity can be seen as a consequence of the von Neumann projection postulate, or of the entanglement of the system with the measurement apparatus (for instance, within dBB theory, the Bohmian positions within this apparatus play an important role [14]).
@@ 2.2 Different views in different Lorentzian frames
We now consider a double interference experiment which, in the Lorentz frame $\mathcal{L}_{0}$ of the laboratory, corresponds to the scheme shown in figure 1. In another Lorentz frame, the same experiment is seen differently. For instance, figure 4 shows a time-space diagram describing the experiment in a frame $\mathcal{L}_{1}$ moving with a positive velocity $v$ along axis $Ox$. The output beam splitter is then crossed in interferometer $\mathcal{A}$ before than in interferometer $\mathcal{B}$. Thus, in this Lorentz frame, it makes sense to consider a simultaneous measurement of the output channel in interferometer $\mathcal{A}$ while the other particle is still in interferometer $\mathcal{B}$; the probability of any result can then be obtained by time integration of the current in this frame. To obtain this probability, we consider the same measurement performed in the laboratory frame $\mathcal{L}_{0}$, which becomes a time-shifted measurement experiment and corresponds to the configuration of figure 2. The probability of this two time measurement in the laboratory frame is then obtained by using the standard Schrödinger theory and the expression (4a) of the state vector. Finally, we use the fact that probabilities are relativistic scalars to obtain the probability of the measurement as seen in the moving frame $\mathcal{L}_{1}$; from this infer properties of the probability current and bi-trajectories in this frame.
For the same reason as in § 1, if the detectors are inserted in channel $A_{2}$ and $b_{1}$, the probability vanishes, but does not if they are inserted in channel $A_{2}$ and $b_{2}$. All bi-trajectories that reach $A_{2}$ therefore have to pass through $b_{2}$. If we extend the (unperturbed) bi-trajectories by continuity we see that, in Lorentz frame $\mathcal{L}_{1}$, all bi-trajectories reaching outputs $A_{2}$ and $B_{2}$ necessarily pass through $b_{2}$. Since, as above, expression (3) shows that no bi-trajectory passes through channels $a_{2}$ and $b_{2}$, we conclude that these bi-trajectories necessarily pass through $a_{1}$ and $b_{2}$.
@@
Figure 4: This figure shows the experiment of figure 1 in another Lorentz frame, moving with a positive velocity $v$ in the direction of the vertical $Ox$ axis in the figure. The particles are assumed to have a non-zero mass. In interferometer $\mathcal{A}$, in some sections the propagation is parallel to the $Ox$ axis, which increases the velocity; in interferometer $\mathcal{B}$ it is antiparallel, which decreases the velocity. As a consequence, the crossing of the output beam splitter occurs first in interferometer $\mathcal{A}$ as in figure 2.